OpenStudy (anonymous):

In a poll in a school, students picked whether they liked apples or oranges more. The ratio for this was 5:8. Five absent students were polled the next day and each of them preferred apples and the new ratio was 5:6. How many people were originally polled?

OpenStudy (anonymous):

Say there are x students of which y like apples y:x-y = 5:8 y+5:x-y =5:6

OpenStudy (anonymous):

You arent making any sense. whats the answer then?

OpenStudy (dumbcow):

it might be easier to look at it in fraction form y/(x-y) = 5/8 (y+5)/(x-y) = 5/6 y is the students who like apples, x is the total num of students x-y students like oranges you have to solve the system of proportions for x and y

OpenStudy (anonymous):

wouldnt the denominator for the second one be (x-y+5)

OpenStudy (dumbcow):

no because the num of students that liked oranges did not change

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so... what next?

OpenStudy (dumbcow):

i would solve one of the equations for y, then substitute into the other equation in order to find x

OpenStudy (anonymous):

what do you get as your answer?

OpenStudy (dumbcow):

x = 39 y = 15

OpenStudy (anonymous):

Thats correct! I'm a bit slow on systems... Thanks for showing me how to do it.